# JMA: a General Algorithm to Craft Nearly Optimal Targeted Adversarial Examples

### v1

@@ 6.3 Multi-label classification
7 Discussion
@@
Our experiments showed that JMA requires a small number of iterations to find an adversarial example, much smaller than the other algorithms, and gets an ASR that is generally higher (in some cases, much higher) for a similar distortion. It is worth stressing that the number of updates $\bar{n}_{it}$ reported in the tables takes into account also the steps of the binary search carried out at the end of the iterations (see Section 5.4). Specifically, $it=(v-1)+n_{bs}$, where $v$ is the number of ’for’ loops, that is, the number of image updates (see Algorithm 1), and $n_{bs}$ is set equal to $6$. Then, the $\bar{n}_{it}$ values reported in the various tables reveal that JMA can indeed find an adversarial image in one shot in many cases, see for instance Table IV for CIFAR-10 ECOC classification and Table VI for the multi-label classification where $\bar{n}_{it}$ is lower than 7 for the attack settings with larger $\epsilon$ [^10] Note that, even when a one-shot attack is possible, a perturbation applied to the image with a weak strength $\epsilon$ may not result in an adversarial image, thus requiring that more iterations are run.. This confirms a-posteriori the validity of one of the most basic assumptions underlying our approach, namely the local linear behavior, that, when it holds, allows to perform the attack in one-shot. When the linear approximation does not hold in the close vicinity of the initial point, and possibly of other points obtained after the updating, the perturbation is added with a small $\epsilon$ in order to change the initial point, and the Jacobian is recomputed (see Algorithm 1). By inspecting the tables (see for instance Table II), we can see that, when this happens, it is often preferable to use a not-too-small $\epsilon$ so that the point is moved farther from the initial point, where the behavior of the local function can be closer to linear, instead of remaining in the vicinity of the initial point where the local approximation may still be not satisfied. Doing so, a high ASR can be achieved with a reduced number of iterations and a similar PSNR.
8 Conclusions
@@
We have proposed a general, nearly optimal, targeted attack, that can solve the original formulation of the adversarial attack by Szegedy et al. under a first order approximation of the network function. The method resorts to the minimization of a Mahalanobis distance term, which depends on the Jacobian matrix taking into account the effort necessary to move the output representation of the image in a given direction. By exploiting the Wolfe duality, the minimization problem is reduced to a non-negative least square (NNLS) problem, that can be solved numerically. The experiments show that the JMA attack is effective against a wide variety of DNNs adopting different output encoding schemes, including networks using error correction output encoding (ECOC) and multi-label classification, outperforming existing attacks in terms of larger attack success rate, lower distortion and lower complexity. In particular, in the multi-label case, our experiments show that JMA is capable to simultaneously change up to 10 (random) labels of the output label vector for VOC2012 classification with a very high ASR (and, in some cases, it can also change all the 20 output labels), a capability that is completely out of reach of the attacks proposed so far. JMA remains effective also in the case of one-hot encoding, where it has much reduced computational complexity with respect to the C&W attack.
Future work will focus on exploiting the JMA attack in a black-box attack scenario, to develop powerful targeted attacks with certain transferability properties. Exploiting the proposed attack for adversarial training, to increase robustness of general classification models against targeted attacks, is also another interesting research direction.
Acknowledgment
This work has been partially supported by the Italian Ministry of University and Research under contract PRIN 2017 2017Z595XS-001 (PREMIER project).

### v2

@@ VII-C Multi-label classification (VOC2012 - InceptionV3)
TABLE X: Results for VOC2012 using different architectures (target label vector from the training set): (a) ResNet50, (b) ViT-B/16 and (c) CLIP+MLP.
Parameters ASR bASR MSE SSIM $\bar{n}_{it}$ Time(s)
JMA (0.05, 200) 1.00 1.0000 3.3e-6 0.999 32.50 2.88
(0.1, 200) 1.00 1.0000 3.8e-6 0.999 21.96 1.60
(0.5, 200) 1.00 1.0000 1.4e-5 0.993 12.84 0.79
(1, 200) 1.00 1.0000 4.0e-5 0.982 11.84 0.80
ML-C&W (0.01,1000) 0.58 0.9583 3.1e-5 0.989 761.00 307.27
LOTS 2000 0.83 0.9610 2.0e-2 0.386 34.84 138.20
MLA-LP 100 0.22 0.9500 3.5e-4 0.964 10.98 648.39
SemA-ML 300 0.31 0.9167 9.3e-5 0.969 5.30 5.45
(a) ResNet50
Parameters ASR bASR MSE SSIM $\bar{n}_{it}$ Time(s)
JMA (0.05, 200) 0.99 0.9993 4.4e-5 0.990 26.01 3.89
(0.1, 200) 1.00 1.0000 4.4e-5 0.990 17.65 2.22
(0.5, 200) 1.00 1.0000 3.9e-5 0.991 9.49 0.85
(1, 200) 1.00 1.0000 4.6e-5 0.989 8.48 0.69
ML-C&W (0.01, 1000) 0.81 0.9825 3.3e-5 0.992 635.34 996.74
LOTS 2000 0.84 0.9750 1.5e-2 0.603 38.00 217.00
MLA-LP 100 0.52 0.9705 6.0e-4 0.910 5.79 787.87
SemA-ML 300 0.16 0.8835 8.6e-5 0.974 17.91 11.52
(b) ViT-B-16
Parameters ASR bASR MSE SSIM $\bar{n}_{it}$ Time(s)
JMA (0.05, 200) 1.00 1.0000 1.6e-6 0.999 28.65 4.46
(0.1, 200) 1.00 1.0000 2.3e-6 0.999 19.72 2.81
(0.5, 200) 1.00 1.0000 1.2e-5 0.994 11.88 1.30
(1, 200) 1.00 1.0000 3.5e-5 0.984 11.07 1.13
ML-C&W (0.01, 1000) 0.14 0.9095 4.8E-05 0.983 684.75 1613.82
LOTS 2000 0.97 0.9818 1.3E-02 0.496 19.59 106.17
MLA-LP 100 0.11 0.9400 2.1E-04 0.977 5.57 710.85
SemA-ML 300 0.09 0.8788 9.9E-05 0.973 1.00 21.84
(c) CLIP+MLP
TABLE XI: Results for VOC2012, for a 10-bit flipping attack (10 bits out of 20 are flipped). (a) ResNet50, (b) ViT-B/16 and (c) CLIP+MLP.
Parameters ASR bASR MSE SSIM $\bar{n}_{it}$ Time(s)
JMA (0.05, 200) 0.00 0.5483 NA NA NA NA
(0.1, 500) 0.94 0.9780 4.8e-5 0.984 291.08 25.76
(0.5, 500) 1.00 1.0000 2.4e-4 0.916 166.60 14.66
ML-C&W (0.01, 1000) 0.37 0.8805 5.0e-4 0.884 631.78 275.72
(0.5, 2000) 0.59 0.8808 1.1e-1 0.155 1108.74 1154.25
(0.1, 3000) 0.42 0.8710 1.6e-2 0.464 581.57 1747.41
(0.1, 5000) 0.45 0.8690 1.6e-2 0.470 756.03 3340.60
MLA-LP 100 0.00 0.5625 NA NA NA NA
SemA-ML 300 0.00 0.4980 NA NA NA NA
(a) ResNet50
Parameters ASR bASR MSE SSIM $\bar{n}_{it}$ Time(s)
JMA (0.05, 200) 0.65 0.9073 1.1e-4 0.974 154.64 25.00
(0.1, 500) 1.00 1.0000 1.6e-4 0.967 126.64 19.87
(0.5, 500) 1.00 1.0000 2.6e-4 0.948 63.89 9.79
ML-C&W (0.01, 1000) 0.51 0.8680 5.5e-4 0.903 582.63 996.45
(0.5, 2000) 0.74 0.8888 5.6e-2 0.272 358.84 3230.31
(0.1, 3000) 0.62 0.8445 1.2e-2 0.584 912.28 4074.83
(0.1, 5000) 0.66 0.8445 1.9e-2 0.565 1497.76 6674.43
MLA-LP 100 0.00 0.5635 NA NA NA NA
SemA-ML 300 0.00 0.4995 NA NA NA NA
(b) ViT-B/16
Parameters ASR bASR MSE SSIM $\bar{n}_{it}$ Time(s)
JMA (0.05, 200) 1.00 0.9975 3.8e-6 0.998 59.16 8.70
(0.1, 500) 1.00 1.0000 5.8e-6 0.997 41.45 5.86
(0.5, 500) 1.00 1.0000 3.3e-5 0.986 22.73 2.79
ML-C&W (0.01, 1000) 0.51 0.8680 5.5e-4 0.903 582.63 2052.44
(0.5, 2000) 0.74 0.8888 5.6e-2 0.272 358.84 3600.40
(0.1, 3000) 0.62 0.8445 1.2e-2 0.584 912.28 4149.43
(0.1, 5000) 0.66 0.8445 1.9e-2 0.565 1497.76 4867.13
MLA-LP 100 0.00 0.5755 NA NA NA NA
SemA-ML 300 0.00 0.4970 NA NA NA NA
(c) CLIP+MLP
VII-D Multi-label classification (VOC2012 - ResNet50, ViT-B/16, CLIP+MLP)
To validate the effectiveness of JMA against different types of architectures, we also trained a residual network (ResNet50), a Transformer-based architecture (ViT-B/16) and a network based on pre-trained vision-language models (CLIP), obtained by adding a lightweight classification head - namely a Multi-Layer Perceptron - on top of CLIP features (CLIP+MLP). The models were trained using the multi-label cross-entropy loss with Adam optimizer with lr = $10^{-3}$, batch size of 32 samples, and early stopping based on the mAP computed on the validation set. The ResNet50, ViT-B/16, and CLIP+MLP models obtained a mAP score on the test set of 0.94, 0.87, and 0.90, respectively.
Table X and XI report the results respectively in the case of attack target vector chosen randomly from the training set and in the case where 10 bit at random are modified in the label vector. For a more comprehensive evaluation on the quality of attacked images and the amount of distortion introduced by the attack, in the tables we also report the SSIM in addition to MSE. We see that the state-of-the-art methods, especially ML-C&W and MLA-LP, achieve worse performance than in the InceptionV3 case, while LOTS is more effective. With regard to JMA, performance are very good and ASR = 1 can be obtained in all the cases and in both attack settings. Furthermore, JMA is the attack introducing the lowest distortion, with an MSE in the range $10^{-6},10^{-4}]$. The SSIM is also very good, being always above 0.99 in the first attack setting, and remaining good also in the challenging 10-bit flips attack. Compared to the InceptionV3 case, JMA takes more iterations to run ($\bar{n}_{it}$ is larger), however the algorithm is extremely fast and the attacked image can be found in a few seconds in most cases.
TABLE XII: Results of the 10-bit flipping attack on MS-COCO dataset.
Parameters ASR bASR MSE SSIM $\bar{n}_{it}$ Time(s)
JMA (0.05,200) 1.00 1.0000 7.7e-6 0.997 53.80 76.86
(0.1, 500) 1.00 1.0000 9.5e-6 0.997 38.59 51.66
(0.5, 500) 1.00 1.0000 4.4e-5 0.984 30.25 40.24
ML-C&W (0.01, 1000) 0.00 0.7649 NA NA NA NA
(0.5, 2000) 0.00 0.7984 NA NA NA NA
(0.1, 3000) 0.00 0.7578 NA NA NA NA
MLA-LP 100 0.00 0.8892 NA NA NA NA
SemA-ML 300 0.00 0.8518 NA NA NA NA
(a) Swin-B
Parameters ASR bASR MSE SSIM $\bar{n}_{it}$ Time(s)
JMA (0.05, 200) 0.46 0.9177 8.9e-6 0.996 149.17 98.73
(0.1, 500) 1.00 1.000 1.9e-5 0.993 163.08 109.78
(0.5, 500) 1.00 1.000 1.4e-4 0.957 101.66 66.38
ML-C&W (0.01, 1000) 0.21 0.9373 5.3e-4 0.905 320.93 1377.91
(0.5, 2000) 0.11 0.9476 1.1e-1 0.222 853.68 3154.67
(0.1, 3000) 0.19 0.9331 2.1e-2 0.481 793.53 4507.12
MLA-LP 100 0.00 0.8816 NA NA NA NA
SemA-ML 300 0.00 0.8596 NA NA NA NA
(b) CLIP+MLP
TABLE XIII: Results of the 10-bit flipping attack on NUS-WIDE dataset (the network is CLIP+MLP).
Parameters ASR bASR MSE SSIM $\bar{n}_{it}$ Time(s)
JMA (0.05,200) 0.99 0.9975 5.1e-6 0.998 81.65 48.80
(0.1, 500) 1.00 1.00 7.8e-6 0.996 59.96 35.23
(0.5, 500) 1.00 1.00 5.3e-5 0.976 36.53 20.35
ML-C&W (0.01, 1000) 0.00 0.7559 NA NA NA NA
(0.5, 2000) 0.00 0.7612 NA NA NA NA
(0.1, 3000) 0.00 0.7070 NA NA NA NA
MLA-LP 100 0.00 0.8843 NA NA NA NA
SemA-ML 300 0.00 0.8630 NA NA NA NA
TABLE XIV: Comparison between JMA and SemA-ML [11] on the NUS-WIDE dataset in the case of semantic consistent multi-label attacks.
Parameters ASR bASR MSE SSIM $\bar{n}_{it}$ Time(s)
JMA (0.05, 200) 1.00 1.0000 1.2e-6 0.999 26.98 14.16
(0.1, 500) 1.00 1.0000 1.8e-6 0.999 19.57 9.39
(0.5, 500) 1.00 1.0000 1.2e-5 0.994 12.33 4.76
SemA-ML 300 0.20 0.9568 7.2e-5 0.978 36.43 5.10
(a) Target label vector from training dataset. The average number of bits that the attack is required to flip is 4.37 bit.
Parameters ASR bASR MSE SSIM $\bar{n}_{it}$ Time(s)
JMA (0.05, 200) 1.00 1.0000 6.6e-7 1.000 19.03 8.84
(0.1, 500) 1.00 1.0000 1.0e-6 0.999 14.17 5.77
(0.5, 500) 1.00 1.0000 6.6e-6 0.996 9.58 2.88
SemA-ML 300 0.48 0.9709 5.8e-5 0.975 13.65 4.20
(b) Attack removing all labels/bits. The average number of bits that the attack is required to flip is 2.90 bit.
VII-E Multi-label classification (MS-COCO, NUS-WIDE)
To validate the effectiveness of JMA against more challenging datasets, we considered the MS-COCO [41] and NUS-WIDE [42] datasets, containing 80 and 81 labels respectively, resulting in a huge number of possible label vectors. For the MS-COCO dataset, we trained a Swin-B architecture (which obtained a better mAP than the ViT-B/16) and CLIP+MLP, with input size of $224\times 224$ pixels. The models were trained on the original split of the dataset using the Asymmetric Loss (ASL) proposed in [43], with Adam optimizer, lr = $10^{-3}$ and batch size of 32 samples. Early stopping on the validation set was employed based on mAP. The Swin-B and CLIP+MLP models scored an mAP of about 0.82 and 0.81, which is aligned with benchmarks for similar input shapes. For the NUS-WIDE case, we split the dataset into 129,431 training, 32,358 validation, and 107,759 test images. A CLIP+MLP architecture was trained, using the same learning setting as for MS-COCO, obtaining a mAP score of 0.59, which is aligned with banchmarks.
Table XII reports the results on MS-COCO in the 10-bit flipping attack scenario. We see that the architecture has a noticeable impact on the performance of the attacks. For the state-of-the-art methods, attacking the Swin-B network is more difficult. Specifically, the bASR values achieved by ML-C&W range from 0.75 to 0.80 and from 0.93 to 0.95 across the various parameter settings respectively in the Swin-B and CLIP+MLP case, thus being similar to the VOC2012 case. However, the ASR is much lower and only very few images can be fully attacked in the CLIP+MLP case, while in the case of Swin-B the ASR is 0. On the contrary, MLA-LP and SemA-ML work better on MS-COCO w.r.t. the VOC2012 case and the bASR is higher. Regarding JMA, in both cases the attack can reach ASR = 1 with low attack distortion, thus surpassing the state-of-the-art. In the CLIP+MLP case, the attack takes more time to run and the distortion introduced by the attack is larger, yet the SSIM remains above 0.99.
For the NUS-WIDE dataset, the results of the 10-bit flipping attack are reported in Table XIII. We see that the ASR is always 0 for the state-of-the-art methods, while attacking with JMA resulted in ASR = 1. Results in terms of distortion and computational efficiency are similar to the best results on MS-COCO, achieved using Swin-B.
It is worth observing that SemA-ML is a multi-label attack which enforces that semantic consistency among the labels is retained by the attack. Differently from the case of VOC2012 and MS-COCO, in the NUS-WIDE dataset, labels correspond to both categories and subcategories (e.g. , ‘vehicle’ and ‘car’ are distinct labels). Hence, semantic consistence may not be verified when bits are flipped at random (e.g., the label ‘car’ may be turned on while the label ‘vehicle’ remains off, leading to a semantically inconsistent vector). Hence, to compare JMA with SemA-ML in a setting that satisfies this method’s working assumptions, we also ran some tests considering the following simple attack settings, where semantic consistency is satisfied: i) target label vector chosen at random from the training/test set; ii) target vector with all labels/bits set to zero. The results in these cases are reported in Table XIV. SemA-ML can now attack most of the labels and reach the target in some cases (and the ASR is no longer 0). This is especially the case when the attack is asked to set all the bits to zero (in this case, the attack involves flipping a lower number of bits - 2.90 on the average). In this case, SemA-ML can achieve bASR = 0.971 and ASR = 0.48. In any case, the performance of JMA remains largely superior, achieving ASR = 1 in all the settings with a lower MSE and an SSIM above 0.99 (still with a lower complexity).
VII-F JMA as a one-step attack
@@
Experiments show that JMA requires a small number of iterations to find an adversarial example, much smaller than the other algorithms, and gets an ASR that is generally higher (in some cases, much higher) for a similar distortion. It is worth stressing that the number of updates $\bar{n}_{it}$ reported in the tables takes into account also the steps of the binary search carried out at the end of the iterations (see Section VI-D). Specifically, $it=(v-1)+n_{bs}$, where $v$ is the number of ‘for’ loops, that is, the number of image updates (see Algorithm 1), and $n_{bs}$ is set equal to $6$. Therefore, JMA can indeed find an adversarial image in one shot in a considerable number of cases, see for instance Table IV for CIFAR-10 ECOC classification and Table 11 for the multi-label classification, where $\bar{n}_{it}$ is lower than 7 for the attack settings with larger $\epsilon$ [^12] Note that, even when a one-shot attack is possible, a perturbation applied to the image with a weak strength $\epsilon$ may not result in an adversarial image, thus requiring that more iterations are run.. This confirms a-posteriori the validity of one of the most basic assumptions underlying our approach, namely the local linear behavior, that, when it holds, allows to perform the attack in one-shot. When the linear approximation does not hold in the close vicinity of the initial point, and possibly of other points during the updates, the perturbation is added with a small $\epsilon$ in order to change the initial point, and the Jacobian is recomputed (see Algorithm 1). By inspecting the tables (see for instance Table I), we can see that, when this happens, it is often preferable to use a not-too-small $\epsilon$ so that the point is moved farther from the initial point, where the behavior of the local function can be closer to linear, instead of remaining in the vicinity of the initial point where the local approximation may still be not satisfied. Doing so, a high ASR can be achieved with a reduced number of iterations and a similar distortion.
VIII Concluding remarks
@@
We have proposed a general, nearly optimal, targeted attack, that can solve the original formulation of the adversarial attack by Szegedy et al. under a first order approximation of the network function. The method resorts to the minimization of a Mahalanobis distance term, which depends on the Jacobian matrix taking into account the effort necessary to move the feature representation of the image in a given direction. By exploiting the Wolfe duality, the minimization problem is reduced to a non-negative least square (NNLS) problem, that can be solved numerically. The experiments show that the JMA attack is effective against a wide variety of DNNs adopting different output encoding schemes, including networks using error correction output encoding (ECOC) and in particular multi-label classification, outperforming existing attacks in terms of higher ASR, lower distortion and lower complexity, with attack capabilities far exceeding those of existing attacks. JMA remains effective also in the case of one-hot encoding, with much reduced computational complexity with respect for instance to the C&W attack.
Finally, we observe that JMA can inspire new defences. In particular, given its efficiency, JMA can be used to reduce the load of adversarial training and develop effective adversarial training defences [2, 44, 45]. Furthermore, while with the exception of scattered works ([44]) adversarial training is primarly limited to applications of single-label classification, JMA can be used to implement effective adversarial training also for other types of classifiers (e.g. multi-label classifiers). In addition, thanks to its efficiency, JMA could also be used to perform adversarial training of provably robust multi-label classifiers exploiting randomized smoothing [46]. In fact, similarly to what has been done for standard single-label randomized smoothing classifiers ([47]), by training with JMA attacked samples it could be possible to boost the provable robustness of smoothed multi-label classifiers. Future work could also focus on the extension of JMA to a black-box attack scenario, to develop powerful targeted attacks with certain transferability properties.